Note: Answer the Question(s) by Using the TI-30XA (or other approved) Calculator. IMPORTANT: Make sure you carry all decimals until you reach your answer, then round to four decimals Use the two-way table below to answer the following question Kids Cars Total 10 15 Total 30 Question 1 P(<3 cars | 3 to 5 kids)- Question 2 P(>3 kids or 2 cars)- Click to Grade Your Work

PLEASE CARRY ALL DECIMALS UNTIL FINAL ANSWER!!
Answer #1

Question 1

Here we want to find:

P( < 3 cars | 3 to 5 Kids) = P( < 3 cars and 3 to 5 kids ) / P( 3 to 5 Kids) = (13/30)/(15/30) = 13/15 = 0.8667

Question 2

P( > 2 Kids or >= 2 cars ) = P(> 2 Kids) + P(>= 2 cars) - P(> 2 Kids and >= 2 cars) = (20/30) + (20/30) - (12/30)

= (20+20-12)/30 = 28/30 = 14/15 = 0.9333